# REALM: A Dataset of Real-World LLM Use Cases

### v1

@@ Abstract
Large Language Models (LLMs), such as the GPT series, have driven significant industrial applications, leading to economic and societal transformations. However, a comprehensive understanding of their real-world applications remains limited. To address this, we introduce REALM, a dataset of over 94,000 LLM use cases collected from Reddit and news articles. REALM captures two key dimensions: the diverse applications of LLMs and the demographics of their users. It categorizes LLM applications and explores how users’ occupations relate to the types of applications they use. By integrating real-world data, REALM offers insights into LLM adoption across different domains, providing a foundation for future research on their evolving societal roles. A dedicated dashboard https://realm-e7682.web.app/ presents the data.
@@
Jingwen Cheng1, Kshitish Ghate1, Wenyue Hua2, William Yang Wang2, Hong Shen1, Fei Fang1 1Carnegie Mellon University, Pittsburgh, USA 2University of California, Santa Barbara, USA {chengjw21, kghate, hongs, feif}@andrew.cmu.edu {wenyuehua, william}@cs.ucsb.edu
@@ 1 Introduction
Currently, there remains limited understanding of how LLM are deployed and utilized in real-world scenarios despite widespread adoption. Existing studies generally fall into two primary categories. The first uses occupation-activity typologies to measure LLM’s influence on various tasks, aiming to assess task-exposure rates and impact on occupations (Eloundou et al., 2023; Felten et al., 2018; Brynjolfsson et al., 2018; Webb, 2020). However, these studies often lack empirical grounding, weakening their validity. The second category analyzes large datasets from social media and other public sources to evaluate sentiment and trends around LLM applications (Leiter et al., 2023; Miyazaki et al., 2024; Koonchanok et al., 2024; Naing and Udomwong, 2024), but these studies are hindered by a lack of comprehensive taxonomies and rely on simplistic methods, such as keyword filtering, leading to low precision and recall.
@@ 2 Related Work
In contrast to structured taxonomy-based approaches, empirical research leveragesNaing and Udomwong (2024) social media data to explore real-world applications of LLMs. For instance, Leiter et al. (2023) analyzes sentiment trends and employs topic modeling to investigate the impacts of LLMs across domains such as education and technology, but utilizing a highly simplified taxonomy. Miyazaki et al. (2024) and Koonchanok et al. (2024) incorporate occupation-level analyses based on O*NET and Indeed to examine how different professions perceive and interact with LLMs. Miyazaki et al. (2024) further analyzes user-provided prompts and applies topic modeling to understand user interactions with LLMs. However, these studies often depend on keyword filtering or hashtag-based data collection methods, which are imprecise and fail to provide a comprehensive taxonomy due to the necessity of omitting many keywords that may have multiple meanings. Beyond social media, other research has examined LLM adoption through alternative data sources with varying focuses including scientific literature Gao and Wang (2024) and large-scale user interaction logs with LLMs Zheng et al. (2023).
Building upon the aforementioned works, REALMintroduce a taxonomy-driven approach coupled with a semantically rich filtering process. We systematically link LLM applications to specific occupations and use cases using both social media and news platforms with high-performing data collection and cleaning pipeline. This enables a precise and nuanced understanding of how different professional sectors are integrating LLMs into their workflows, offering a robust foundation for future research.
@@ 3 Dataset Construction
The data collection process begins with keyword-based extraction to isolate LLM-related content, using a comprehensive list of model names obtained from GitHub [^3] https://github.com/Hannibal046/Awesome-LLM. Then we perform essential data cleaning to ensure the quality of the collected content. This includes removing duplicates, unusable tweets, stop words, and non-textual elements, ensuring the data is well-prepared for further processing. We employ a fine-tuned RoBERTa model to filter irrevelant data, ensuring relevance by identifying descriptions of LLM use cases or potential applications. This filtering process prioritizes recall over precision to include all pertinent data, achieving high recall scores of 0.94 for news articles and 0.95 for Reddit posts after human validation. Following the initial filtering, a four-stage annotation pipeline, detailed in Section 3, is applied to further refine the dataset by enhancing precision.
@@
To evaluate the performance of the whole pipeline, two expert annotators labeled a total of 3,000 data points. See Appendix A.3 for more information.
@@ 4 Analysis of REALM Dataset
Figures 2(a) illustrate the number of LLM use cases categorized by application type within News and Reddit sources, respectively. These figures reveal that Content Creation and Content Synthesis are the most prevalent LLM use cases, including writing assistance, summarization, and code generation. Additionally, applications related to Decision-making, Digital Assistance, and Process Automation have experienced substantial growth.
@@ 5.2 Conclusion
While REALM provides valuable insights into the utilization of LLMs across various professions and tasks, it is subject to several limitations. Firstly, the dataset is exclusively derived from Reddit and news articles, excluding other significant platforms such as Twitter, Facebook, and LinkedIn. Secondly, the inherent characteristics of the selected sources contribute to skewed distributions within the dataset. Reddit’s user base is predominantly technical, and news articles often emphasize business-related topics, leading to an overrepresentation of fields such as Computer Science and Business, while less prominent sectors like Transportation and Farming remain underrepresented. Addressing these biases and establishing accurate denominators for normalization purposes lie beyond the current scope of REALM.
@@ A.1 Taxonomy Selection
A.2 Details of the Four-Module Annotation Pipeline
@@
A.3 Data Annotation Details
@@
The first task involves annotating 2,000 data points selected from the set of Reddit posts and news articles after keyword filtering. These data points were categorized into “Use Case” and “Non-Use Case” labels to train a RoBERTa model, which acted as an automated classifier, thus optimizing resource use and time efficiency. The IAA for this task reached 0.89, and the resulting RoBERTa model achieved high recall scores, as discussed in Section 3.
@@
A.4 Prompts

### v2

@@ Abstract
Large Language Models (LLMs), such as the GPT series, have driven significant industrial applications, leading to economic and societal transformations. However, a comprehensive understanding of their real-world applications remains limited. To address this, we introduce REALM, a dataset of over 94,000 LLM use cases collected from Reddit and news articles. REALM captures two key dimensions: the diverse applications of LLMs and the demographics of their users. It categorizes LLM applications and explores how users’ occupations relate to the types of applications they use. By integrating real-world data, REALM offers insights into LLM adoption across different domains, providing a foundation for future research on their evolving societal roles. An interactive dashboard (https://realm-e7682.web.app/) is provided for easy exploration of the dataset.
@@
Jingwen Cheng1, Kshitish Ghate1, Wenyue Hua2, William Yang Wang2, Hong Shen1, Fei Fang1 1Carnegie Mellon University, Pittsburgh, USA 2University of California, Santa Barbara, USA {chengjw21}@gmail.com {kghate, hongs, feif}@andrew.cmu.edu {wenyuehua, william}@cs.ucsb.edu
@@ 1 Introduction
Currently, there remains limited understanding of how LLMs are deployed and utilized in real-world scenarios despite widespread adoption. Existing studies generally fall into two primary categories. The first uses occupation-activity typologies to measure LLM’s influence on various tasks, aiming to assess task-exposure rates and impact on occupations (Eloundou et al., 2023; Felten et al., 2018; Brynjolfsson et al., 2018; Webb, 2020). However, these studies often lack empirical grounding, weakening their validity. The second category analyzes large datasets from social media and other public sources to evaluate sentiment and trends around LLM applications (Leiter et al., 2023; Miyazaki et al., 2024; Koonchanok et al., 2024; Naing and Udomwong, 2024), but these studies are hindered by a lack of comprehensive taxonomies and rely on simplistic methods, such as keyword filtering, leading to low precision and recall. While some industry research (e.g., from LinkedIn, Anthropic, Goldman Sachs) on LLM impact is available, it is often not publicly accessible or prohibitively costly, limiting its use in academic research.
@@ 2 Related Work
In contrast to structured taxonomy-based approaches, empirical research leverages social media data to explore real-world applications of LLMs Naing and Udomwong (2024). For instance, Leiter et al. (2023) analyzes sentiment trends and employs topic modeling to investigate the impacts of LLMs across domains such as education and technology, but utilizing a highly simplified taxonomy. Miyazaki et al. (2024) and Koonchanok et al. (2024) incorporate occupation-level analyses based on O*NET and Indeed to examine how different professions perceive and interact with LLMs. Miyazaki et al. (2024) further analyzes user-provided prompts and applies topic modeling to understand user interactions with LLMs. However, these studies often depend on keyword filtering or hashtag-based data collection methods, which are imprecise and fail to provide a comprehensive taxonomy due to the necessity of omitting many keywords that may have multiple meanings. Beyond social media, other research has examined LLM adoption through alternative data sources with varying focuses, including scientific literature Gao and Wang (2024), large-scale user interaction logs with LLMs Zheng et al. (2023); Zhao et al. (2024), and proprietary usage data from commercial LLM deployments Handa et al. (2025). The latter offers a comprehensive analysis of millions of conversations with Claude, revealing correlations with user occupations, salaries, and job zones. However, their data is not publicly available, and the results are limited to a relatively short time frame.
Building upon the aforementioned works, REALM introduce a taxonomy-driven approach coupled with a semantically rich filtering process. We systematically link LLM applications to specific occupations and use cases using both social media and news platforms with high-performing data collection and cleaning pipeline. Unlike prior studies based on proprietary or short-term data, REALM continuously aggregates public content across diverse domains over an extended time span, enabling a precise and nuanced understanding of how different professional sectors are integrating LLMs into their workflows, offering a robust foundation for future research.
@@ 3 Dataset Construction
The data collection process begins with keyword-based extraction to isolate LLM-related content, using a comprehensive list of model names obtained from GitHub [^3] https://github.com/Hannibal046/Awesome-LLM. Then we perform essential data cleaning to ensure the quality of the collected content. This includes removing duplicates, unusable tweets, stop words, and non-textual elements, ensuring the data is well-prepared for further processing. We employ a fine-tuned RoBERTa model to filter irrevelant data, ensuring relevance by identifying descriptions of LLM use cases or potential applications. The details of the RoBERTa fine-tuning process are provided in Appendix A.2. This filtering process prioritizes recall over precision to include all pertinent data, achieving high recall scores of 0.94 for news articles and 0.95 for Reddit posts after human validation. Following the initial filtering, a four-stage annotation pipeline, detailed in Section 3, is applied to further refine the dataset by enhancing precision.
@@
To evaluate the performance of the whole pipeline, two expert annotators labeled a total of 3,000 data points.1 See Appendix A.4 for more information.
@@
Figure 2: Visualization of LLM Mentions Across Usage and end-user Professions
@@ 4 Analysis of REALM Dataset
Figure 2(a) illustrates the number of LLM use cases categorized by application type within News and Reddit sources, respectively. These figures reveal that Content Creation and Content Synthesis are the most prevalent LLM use cases, including writing assistance, summarization, and code generation. Additionally, applications related to Decision-making, Digital Assistance, and Process Automation have experienced substantial growth.
@@
Debiasing
To address potential biases arising from the uneven distribution of occupational groups, we conducted a debiasing analysis to normalize raw LLM use case statistics. Certain fields, such as Arts or Business, may dominate the dataset due to higher overall content volume, artificially inflating their apparent LLM adoption.
We addressed this by applying a normalization procedure. We randomly sampled 1,000 data points from both the news and Reddit corpora and manually annotated each item with its associated occupational group. Based on these annotations, we estimated the discussion distribution across occupations: for each group $i$, we computed its discussion proportion $D_{i}$, defined as the fraction of sampled items belonging to group $i$. This captures how prevalent each group is in the overall corpus, independent of LLM-specific content.
Next, we normalized the number of LLM use cases associated with each occupational group by dividing by $D_{i}$, resulting in what we refer to as the LLM exposure rate:
$\text{ExposureRate}_{i}=\frac{\text{NumOfUseCases}_{i}}{D_{i}}$
This metric reflects the relative concentration of LLM-related content within each group, adjusted for how frequently the group appears in the dataset overall. By using this normalization, we aim to distinguish true differences in LLM adoption from artifacts of dataset imbalance.
After debiasing (Figure 3), the exposure rates of previously dominant fields such as Business, Management, and Computer Science decrease markedly relative to their raw use case counts (Figure 3(b)). This indicates that their initial prominence was primarily driven by the sheer volume of content associated with these occupations, rather than particularly high LLM adoption. In contrast, domains such as Education, Healthcare, and Art and Media retain relatively high exposure rates even after normalization, suggesting that LLM-related use cases constitute a larger share of discourse in these fields. Furthermore, the side-by-side comparison in Figure 3(c) highlights notable differences between Reddit and News data, where user-driven domains like Art and Media and Education exhibit higher Reddit exposure, while business-related fields are more prominent in News coverage.
(a) Exposure Rate of News Data.
(b) Exposure Rate of Reddit Data.
(c) Comparison of Exposure Rate between News and Reddit Data.
Figure 3: Visualization of Exposure Rate after debiasing.
@@ 5.2 Conclusion
While REALM provides valuable insights into the utilization of LLMs across various professions and tasks, it is subject to several limitations. Firstly, the dataset is exclusively derived from Reddit and news articles, excluding other significant platforms such as Twitter, Facebook, and LinkedIn. Secondly, while we attempted to mitigate occupational group imbalance through a debiasing analysis (Section 4), our method remains a simplification. Future work could explore more nuanced normalization strategies, such as incorporating engagement signals or semantic relevance.
@@ A.1 Taxonomy Selection
A.2 RoBERTa Fine-tuning Details
Model and Hyperparameters.
We fine-tuned the roberta-base model using the following hyperparameters:
  • Learning Rate: $2\times 10^{-5}$
  • Batch Size: 16
  • Number of Epochs: 10 (applied separately to the two stages of transfer learning)
  • Optimizer: AdamW
  • Class Rebalancing: Applied to mitigate class imbalance, with a weight of 3.0 assigned to the minority class (use cases)
  • Loss Function: CrossEntropyLoss with class weights
  • Maximum Sequence Length: 512
Training Data.
The model was trained on a total of 2,000 annotated data points, comprising 1,000 news articles and 1,000 Reddit posts. Each data point was manually labeled as either a use case or a non-use case. Notably, only approximately 15% of the samples were labeled as use cases, resulting in a significant class imbalance. To address this, we applied class rebalancing techniques during training to ensure effective learning from the underrepresented class.
The training procedure consisted of two phases: an initial joint fine-tuning stage using the combined dataset, followed by separate fine-tuning on the news and Reddit domains to account for their domain-specific characteristics.
Evaluation.
The model was evaluated on a held-out test set of 400 instances, equally split between 200 news articles and 200 Reddit posts. Despite the low prevalence of use cases in the data, the model achieved a high recall score and a relatively high precision score, demonstrating its effectiveness in identifying use cases under imbalanced conditions.
A.3 Details of the Four-Module Annotation Pipeline
@@
A.4 Data Annotation Details
@@
The first task involves annotating 2,000 data points selected from the set of Reddit posts and news articles after keyword filtering. We provide clear guideline on what LLM use cases mean to human annotators, details can be found in Appendix A.5. These data points were categorized into “Use Case” and “Non-Use Case” labels to train a RoBERTa model, which acted as an automated classifier, thus optimizing resource use and time efficiency. The IAA for this task reached 0.89, and the resulting RoBERTa model achieved high recall scores, as discussed in Section 3.
@@
A.5 Annotation Guideline for Use Case Classification
We provided the following guideline on what LLM use cases mean to human annotators (with extensive examples, though not all are listed here):
  • Current Use of LLM: Cases that describe or mention the direct use of large language models (LLMs), including through APIs or products integrated with LLMs that are being used by end-users. These cases always involve clear descriptions of users, or the users themselves are the subjects of the articles, indicating that LLMs or products, software, tools, or systems built on LLMs are being actively used.
    Example: customer service chatbots used by customers, LLM-based writing tools, or developer-facing APIs. Any direct interaction with an LLM to generate content or provide functionality is also categorized here.
  • Potential Use of LLM: Cases that involve the introduction, release, or advertisement of products, tools, or systems built using LLMs, which are mostly speculative or have not yet been adopted by end-users. This excludes ethical discussions or sentiment analysis without actual use.
    Example: announcements of upcoming LLM-powered features or promotional materials for unreleased tools.
We differentiate between current and potential use cases because the two types of posts or articles may vary significantly. Current use cases usually have clearly defined end-users, while potential use cases often lack clear end-user involvement and may simply describe a newly released product. This distinction helps reduce confusion during the annotation process.
A.6 Prompts